# SyncDiff: Synchronized Motion Diffusion for Multi-Body Human-Object Interaction Synthesis

### v2

@@ 5.1 Datasets
To examine our method’s generalizability across various multi-body interaction configurations, we utilize four datasets with different interaction scenarios: TACO [46] (two hands and two objects), CORE4D [99] (two people and one object), GRAB [76] (one or two hands and one object), and OAKINK2 [98] (two hands and one to three objects). We describe data splits for each dataset below.
@@
(3) GRAB: We use an existing data split of unseen subjects from IMoS [22] and that of unseen objects from DiffH2O [13]. Please refer to the two papers for details.
@@
Toward a fair comparison with DiffH2O [13], we follow the setting of DiffH2O for GRAB, where methods are required to generate hand-object manipulations after grasping objects. For the other three datasets, all methods need to synthesize complete multi-body interaction motion sequences without direct guidance.
@@ 5.2 Evaluation Metrics
To evaluate the qualities of synthesized motion sequences comprehensively, we present two types of evaluation metrics focusing on fine-grained physical plausibility and general motion semantics, respectively.
(1) Physics-based metrics measure the physical plausibility of hand-object/human-object interactions and the extent of motion coordination among different bodies. We use the Contact Surface Ratio (CSR) for hand-object settings and the Contact Root Ratio (CRR) for human-object settings to denote the proportion of motion sequences where hand-object/human-object contact occurs. Contact is defined as the hand mesh being within $5$mm of at least one object for CSR, and the two hand roots of a human consistently being within $3$cm of at least one object for CRR. When there are multiple hands or humans, we take the average among all of them. We label the frames based on whether hand-object contact occurs for ground-truth and synthesized motions and then compute their Intersection-over-Union (IoU), denoted as CSIoU. Besides, Interpenetration Volume (IV) and Interpenetration Depth (ID) are incorporated to assess penetration between different bodies.
(2) Motion semantics metrics evaluate high-level motion semantics and its distributions, comprising Recognition Accuracy (RA), Fréchet Inception Distance (FID), Sample Diversity (SD), and Overall Diversity (OD). Following existing evaluations for motion synthesis [40, 39], we train a network to extract motion features and predict action labels using ground-truth motion data. For better feature semantics, the network is trained on the combination of all train, val, and test splits. RA denotes the action recognition accuracy of the network on synthesized motions. FID measures the difference in feature distributions of generated and ground-truth motions. Following  [13], SD represents the mean Euclidean distance between multiple generated wrist trajectories in a single sample, and OD refers to the mean distance between all generated trajectories in a dataset split.
@@ 5.3 Comparison to Existing Methods
Hand-Object Interaction. For hand-object interaction motion synthesis, we compare our SyncDiff to two state-of-the-art approaches MACS [68] and DiffH2O [13]. For MACS, we first generate the motions of all objects in their object trajectory synthesis phase, and then directly use their hand motion synthesis phase to get overall results. For DiffH2O, we use the version without grasp reference input.
@@
Figure 3: Qualitative results from GRAB [76] dataset.
As is shown in Table 1, 2, and 4, our method outperforms MACS and DiffH2O by a large margin with better CSIoU, CSR, IV, and ID. The reason is that our method features more robust alignment and synchronization mechanisms to ensure synchronization which are excluded in existing methods. Figure 2 compares results from different methods qualitatively. In the results from MACS and DiffH2O, the brush often penetrates the plate, and the pose of the brush does not guarantee the bristles being pressed closely against the plate, effectively completing the action. FID and RA further indicate that the motions generated by our method are more realistic. This is still caused by the more precise generation quality of object-object interactions from our method, and our separation of motions at different frequencies ensures that subtle high-frequency periodic movements are not overshadowed, which are crucial for identifying action types. Figure 4 demonstrates the benefits of our synchronization mechanism. Although MACS uses relative representations for some hand-object pairs to ensure firm grasp, due to the lack of synchronization on the complete graphical model, conflicts arise between the left hand and the knife, which should not directly interact. Figure 6 indicates that with a higher demand for coordination quality among objects, our method can also address hand-object and object-object synchronization effectively.
@@
Human-Object Interaction. For human-object interaction synthesis, we compare our method to ComMDM [67], InterGen [42], and OMOMO [40]. ComMDM and InterGen are capable of generating interactions between two people. For OMOMO, we first use their conditional diffusion models to generate object trajectories, and then use their whole pipeline to synthesize complete Multi-body HOI motion.
As shown in Table 3 and Figure 5, our method outperforms all three baseline methods in physics-based and semantics metrics and obtains the most realistic qualities. Generated motions from existing methods suffer from unnatural grasping poses and arm-object interpenetration, while our method mitigates these issues.
@@
Figure 6: Qualitative results on OAKINK2 [98] dataset. The task requires precise contact between objects, where the bottle cap needs to align perfectly with the bottle, and there needs to be a tendency for it to be twisted down in a clockwise spiral.
@@
Table 2: Comparison on GRAB [76] dataset for the post-grasping phase. We conduct experiments on the post-grasping hand-object manipulation phase to compare with DiffH2O [13]. Each column highlights the best method in red, with the second best highlighted in blue. Results of IMoS [22] and DiffH2O are from the original paper of DiffH2O, while MACS [68] results are obtained via our re-implementation.
@@ 5.4 Ablation studies
We examine the effect of our three key designs (frequency-domain motion decomposition, the alignment loss $\mathcal{L}_{\text{align}}$, and the explicit synchronization) separately. The results after removing each of these three components individually are shown as “w/o decompose", “w/o $\mathcal{L}_{\text{align}}$", and “w/o exp sync". Two additional ablations are to remove the two synchronization mechanisms and all three designs, denoted as “w/o $\mathcal{L}_{\text{align}}$, exp sync" and “w/o all", respectively. Results in Table 1, 3, and 4 show that removing any of the above three components can lead to varying extents of performance decline. Compared to the method without any key design, simply incorporating frequency-domain decomposition obtains coherent performance drops under nearly all physics-based metrics due to more challenging multi-body synchronization. Nevertheless, the two synchronization strategies consistently enhance the effects and finally achieve the best results on nearly all evaluation aspects.
Removal of Decomposition. As shown in Figure 7, after removing the decomposition mechanism, once periodic relative motion is involved in the interactions between objects, It becomes easier for the motions to be overly smooth, which is intuitively manifested as two objects being in an almost relative stationary state, making it difficult to complete the action. This also results in poor performance in semantics metrics FID and RA from Table 1,  3, and  4.
@@
Figure 7: Qualitative results from TACO [46] dataset. Periodic relative motion is required between two objects. The color changes from deep to light, representing time passage. It can be observed that after removing the decomposition mechanism, the spatula tends to get stuck in a small area on the plate’s surface.
@@
Removal of $\mathcal{L}_{\text{align}}$ or Explicit Synchronization. As is shown in Figure 5 and 6, removing either one of $\mathcal{L}_{\text{align}}$ or explicit synchronization leads to unreasonable penetration or contact loss between objects or between humans and objects, with explicit synchronization playing a more significant role. This phenomenon is also revealed in quantitative evaluation results from Table  1,  3 and  4.
6 Conclusions
We present SyncDiff, a universal framework for synchronized motion synthesis of multi-body HOI interaction by estimating both data sample scores and alignment scores, and jointly optimizing sample and alignment likelihoods in inference. We also introduce a frequency-domain decomposition to better capture high-frequency, small-amplitude motion details. Experiments on four datasets demonstrate that SyncDiff can be adapted to synthesize synchronized multi-body human-object interaction sequences with any number of humans, hands, and rigid objects. Comparative experiments also demonstrate that in each specific setting, our method achieves better physical accuracy and action semantics than a range of state-of-the-art works. Some limitations of our work and their potential solutions are discussed in the appendix.
@@ A.1 Diffusion Model Basics
Diffusion models [70] and its variants [73, 28], especially latent diffusion, have been widely applied in different tasks, such as high-resolution image generation [63] or video generation [30]. They simulate the data distribution by introducing a series of variables $\{x_{i}\}_{i=1}^{\mathcal{T}}$ with different noise levels. The forward noise process can be represented as
@@
For the inference process, from $\mathcal{T}$ that is large enough so that $p(x_{\mathcal{T}})\approx\mathcal{N}(0,I)$, do reverse sampling step by step, following
@@ A.2 Complete Proof of Our Two Synchronization Mechanisms
First let’s derive some commonly used formulas in diffusion models that we will need in our proof, which the readers might not be familiar with. If you are familiar with the step-by-step denoising formula of diffusion, you can skip directly to Eq  8 and  9.
@@
In general, diffusion models predict the noise $\epsilon$, but in our setting, we need to define the alignment loss later. Thus, it’s more convenient to predict $\hat{x}_{0}$, which is the result after complete denoising. Note that
@@
By removing the coefficient, it becomes our simple reconstruction loss. Of course, in our implementation, we split it into $\mathcal{L}_{\text{DC}}$ and $\mathcal{L}_{\text{AC}}$ for separate supervision. Similarly, in order to derive the alignment loss, our approach is still to express the loss function as some form of negative log-likelihood. Consider a triplet $(\hat{x}_{u,\textbf{b}_{1}},\hat{x}_{u,\textbf{b}_{2}},\hat{x}_{b,\textbf{b}_{2}\rightarrow\textbf{b}_{1}})$, where $\textbf{b}_{1}$, $\textbf{b}_{2}$ are two different bodies. We know that the definition of the score function is $\nabla\log p(x)$, which is just the negative gradient of the negative log-likelihood. As stated in the introduction part, we hope that this score function can guide $\hat{x}_{b,\textbf{b}_{2}\rightarrow\textbf{b}_{1}}$ towards $\text{rel}(\hat{x}_{u,\textbf{b}_{1}},\hat{x}_{u,\textbf{b}_{2}})$. A simple solution is to let $\hat{x}_{b,\textbf{b}_{2}\rightarrow\textbf{b}_{1}}$ follow a distribution $\mathcal{N}(\mu,\sigma^{2})$, where $\mu=\text{rel}(\hat{x}_{u,\textbf{b}_{1}},\hat{x}_{u,\textbf{b}_{2}})$, and $\sigma$ is a parameter that we can tune. Note that a vector $\hat{x}$ satisfying $\hat{x}\sim\mathcal{N}(\hat{\mu},\sigma^{2}I)$ has the probability density function
@@
It is not difficult to obtain that the negative log-likelihood is $\frac{1}{2\sigma^{2}}\|\hat{x}_{b,\textbf{b}_{2}\rightarrow\textbf{b}_{1}}-\text{rel}(\hat{x}_{u,\textbf{b}_{1}},\hat{x}_{u,\textbf{b}_{2}})\|_{2}^{2}$. Summing up for all such pairs $(\textbf{b}_{1},\textbf{b}_{2})$, removing the coefficients, and note that the relative relationships between rigid bodies are bidirectional, we can derive the alignment loss
$\mathcal{L}_{\text{align}}=$ $\sum\limits_{1\leq j_{1}<j_{2}\leq m}2\|\hat{x}_{b,o_{j_{2}}\rightarrow o_{j_{1}}}-\text{rel}(\hat{x}_{u,o_{j_{1}}},\hat{x}_{u,o_{j_{2}}})\|_{2}^{2}$ (13)
$+\sum\limits_{i\in[1,n],j\in[1,m]}\|\hat{x}_{b,h_{i}\rightarrow o_{j}}-\text{rel}(\hat{x}_{u,o_{j}},\hat{x}_{u,h_{i}})\|_{2}^{2}.$
Next, let’s derive the formula for the explicit synchronization mechanism in inference. In our task, during the inference process, suppose we want to derive $\hat{x}^{{}^{\prime}}=\hat{x}_{t-1}$ from $\hat{x}_{t}=\hat{x}$, we first use the denoising backbone to predict the mean value of distribution $\hat{\mu}=\hat{\mu}_{t}$.
In our task, $\hat{x}^{{}^{\prime}}=[\hat{x}^{{}^{\prime}}_{u,o_{1}},\hat{x}^{{}^{\prime}}_{u,o_{2}},\dots,\hat{x}^{{}^{\prime}}_{u,o_{m}},\\ \hat{x}^{{}^{\prime}}_{u,h_{1}},\hat{x}^{{}^{\prime}}_{u,h_{2}},\dots,\hat{x}^{{}^{\prime}}_{u,h_{n}},\hat{x}^{{}^{\prime}}_{b,o_{2}\rightarrow o_{1}},\hat{x}^{{}^{\prime}}_{b,o_{3}\rightarrow o_{1}},\dots,\hat{x}^{{}^{\prime}}_{b,o_{m}\rightarrow o_{m-1}},\\ \hat{x}^{{}^{\prime}}_{b,h_{1}\rightarrow o_{1}},\hat{x}^{{}^{\prime}}_{b,h_{1}\rightarrow o_{2}},\dots,\hat{x}^{{}^{\prime}}_{b,h_{n}\rightarrow o_{m}}]$, and $\hat{\mu}=[\hat{\mu}_{u,o_{1}},\hat{\mu}_{u,o_{2}},\dots,\hat{\mu}_{u,o_{m}},\\ \hat{\mu}_{u,h_{1}},\hat{\mu}_{u,h_{2}},\dots,\hat{\mu}_{u,h_{n}},\hat{\mu}_{b,o_{2}\rightarrow o_{1}},\hat{\mu}_{b,o_{3}\rightarrow o_{1}},\dots,\hat{\mu}_{b,o_{m}\rightarrow o_{m-1}},\\ \hat{\mu}_{b,h_{1}\rightarrow o_{1}},\hat{\mu}_{b,h_{1}\rightarrow o_{2}},\dots,\hat{\mu}_{b,h_{n}\rightarrow o_{m}}]$, so
$\frac{1}{2\sigma^{2}}\|\hat{x}^{{}^{\prime}}-\hat{\mu}\|_{2}^{2}$ $=\frac{1}{2\sigma^{2}}\left(\sum\limits_{i=1}^{n}\|\hat{x}^{{}^{\prime}}_{u,h_{i}}-\hat{\mu}_{u,h_{i}}\|_{2}^{2}\right.$
$\quad+\sum\limits_{j=1}^{m}\|\hat{x}^{{}^{\prime}}_{u,o_{j}}-\hat{\mu}_{u,o_{j}}\|_{2}^{2}$
$\quad+\sum\limits_{1\leq j_{1}<j_{2}\leq m}\|\hat{x}^{{}^{\prime}}_{b,o_{j_{2}}\rightarrow o_{j_{1}}}-\hat{\mu}_{b,o_{j_{2}}\rightarrow o_{j_{1}}}\|_{2}^{2}$
$\quad+\left.\sum\limits_{1\leq i\leq n,1\leq j\leq m}\|\hat{x}^{{}^{\prime}}_{b,h_{i}\rightarrow o_{j}}-\hat{\mu}_{b,h_{i}\rightarrow o_{j}}\|_{2}^{2}\right)$
@@
$\mathcal{P}_{\text{align}}$ $=\sum\limits_{k=1}^{|\mathcal{S}|}\lambda_{k}\|c_{k}-a_{k}\circ b_{k}\|_{2}^{2}$

### v3

@@ 4 Experiments
In this section, we first introduce some basic experimental settings (Section  4.1,  4.2), and then demonstrate the visual and quantitative results in comparison to several baselines (Section  4.3) and ablation settings (Section  4.4). We also provide in-depth analyses for the results.
@@ 4.1 Datasets
To examine our method’s generalizability across various multi-body interaction configurations, we utilize five datasets with different interaction scenarios: TACO [48] (two hands and two objects), CORE4D [109] (two people and one object), GRAB [81] (one or two hands and one object), OAKINK2 [108] (two hands and one to three objects), and BEHAVE [2] (one human and one object). We describe data splits for each dataset below. Detailed dataset statistics and split sizes are shown in Section  D.1.
@@
(3) GRAB: We use an existing data split of unseen subjects from IMoS [24] and that of unseen objects from DiffH2[13]. Please refer to the two papers for details.
@@
(5) BEHAVE: We utilize the official splits of train/test.
Toward a fair comparison with DiffH2[13], we follow the setting of DiffH2O for GRAB, where methods are required to generate hand-object manipulations after grasping objects. For other four datasets, all methods need to synthesize complete multi-body interaction motion sequences.
@@ 4.2 Evaluation Metrics
To evaluate the qualities of synthesized motion sequences comprehensively, we present two types of evaluation metrics focusing on fine-grained contact consistency and general motion semantics, respectively.
(1) Contact-based metrics measure the contact plausibility of hand-object/human-object interactions and the extent of motion coordination among different bodies. We use the Contact Surface Ratio (CSR) for hand-object settings and the Contact Root Ratio (CRR) for human-object settings to denote the proportion of frames where hand-object/human-object contact occurs. Contact is defined as the hand mesh being within $5$mm of at least one object for CSR, and the two hand roots of a human consistently being within $3$cm of at least one object for CRR. When there are multiple hands or humans, we take the average among all of them. We label the frames based on whether hand-object contact occurs for ground-truth or synthesized motions, and then compute their Intersection-over-Union (IoU), denoted as CSIoU. To eliminate potential ambiguities, we’ve provided corresponding pseudocodes in Section  LABEL:C.5. Besides, Interpenetration Volume (IV) and Interpenetration Depth (ID) are incorporated to assess penetration between different bodies.
(2) Motion semantics metrics evaluate high-level motion semantics and its distributions, comprising Recognition Accuracy (RA), Fréchet Inception Distance (FID), Sample Diversity (SD), and Overall Diversity (OD). Following existing evaluations for motion synthesis [42, 41], we train a network to extract motion features and predict action labels using ground-truth motion data. For better feature semantics, the network is trained on the combination of all train, val, and test splits. RA denotes the action recognition accuracy of the network on synthesized motions. FID measures the difference in feature distributions of generated and ground-truth motions. Following DiffH2[13], SD represents the mean Euclidean distance between multiple generated wrist trajectories in a single sample, and OD refers to mean distance between all generated trajectories in a dataset split. To measure semantics from a perceptual aspect, we conduct user studies, comparing our methods with baselines, which is shown in Section  4.5.
@@ 4.3 Comparison to Existing Methods
Hand-Object Interaction. For hand-object interaction motion synthesis, we compare our method to two state-of-the-art approaches, MACS [73] and DiffH2[13]. For MACS, we first generate the motions of all objects in their object trajectory synthesis phase, and then directly use their hand motion synthesis phase to get overall results. For DiffH2O, we use the version without grasp reference input.
Figure 3: Qualitative results from GRAB [81] dataset.
@@
In Figure  3, our method features more realistic contacts and more stable grasping. As is shown in Table 1, 2, and 5, our method outperforms MACS and DiffH2O by a large margin with better CSIoU, CSR, IV, and ID. The reason is that our method features more robust alignment and synchronization mechanisms to ensure synchronization which are excluded in existing methods. In the results from MACS and DiffH2O in Figure  4, the brush often penetrates the plate, and the pose of the brush does not guarantee the bristles being pressed closely against the plate, effectively completing the action. FID and RA further indicate that the motions generated by our method are more semantically realistic. This is caused by the more precise object-object interactions in SyncDiff, and our separation of motions at different frequencies ensures that subtle high-frequency periodic movements are not overshadowed, which are crucial for identifying action types. Figure  5 demonstrates the benefits of our synchronization mechanism. Although MACS uses relative representations for some hand-object pairs to ensure firm grasp, due to the lack of synchronization on the complete graphical model, conflicts arise between the left hand and the knife, which should not directly interact. Figure 6 indicates that with a higher demand for coordination quality among objects, our method can also address hand-object and object-object synchronization effectively.
@@
Figure 6: Qualitative results from OAKINK2 [108] dataset. The task requires precise contact between objects, where the bottle cap needs to align perfectly with the bottle, and there needs to be a tendency for it to be twisted down in a clockwise spiral.
Human-Object Interaction. For human-object interaction synthesis, we compare our method to OMOMO [42] and CG-HOI [20]. For OMOMO, we first use their conditional diffusion models to generate object trajectories, and then use their whole pipeline to synthesize complete Multi-body HOI motion. Cross-attention in CG-HOI is modified for two humans, one object and contact between them.
As shown in Table  3,  4 and Figure 7,  8, our method outperforms CG-HOI and OMOMO in contact-based and semantics metrics and obtains the most realistic qualities. Synthesized motions from existing methods suffer from unnatural grasping poses and arm-object interpenetration, while our method mitigates these issues.
@@
Figure 8: Qualitative results from BEHAVE [2] dataset. Baseline methods suffer from unreasonable grasp poses due to unsynchronized synthesis of body transformations.
@@
Table 2: Comparison on GRAB [81] dataset for the post-grasping phase. Following DiffH2[13], we conduct experiments on the phase where the object has been grasped. Each column highlights the best method in red, with the second best highlighted in blue. Results of IMoS [24] and DiffH2O are from the original paper of DiffH2O, while MACS [73] results are obtained via our re-implementation.
@@ 4.4 Ablation studies
We examine the effect of our three key designs (frequency-domain motion decomposition, the alignment loss $\mathcal{L}_{\text{align}}$, and the explicit synchronization) separately. The results after removing each of these three components individually are shown as “w/o decompose”, “w/o $\mathcal{L}_{\text{align}}$”, and “w/o exp sync”. Two additional ablations are to remove the two synchronization mechanisms and all three designs, denoted as “w/o $\mathcal{L}_{\text{align}}$, exp sync” and “w/o all”, respectively. More details can be found in Section  C.4. Results in Table 1, 3, 5 show that removing any of the three components can lead to varying extents of performance decline.
Removal of Decomposition. As shown in Figure 9, after removing the decomposition mechanism, once periodic relative motion is involved in the interactions between objects, it becomes easier for high-frequency motions to be neglected, which is intuitively shown as two objects being in an almost relative stationary state, making it difficult to complete the action. This also results in poor performance in semantics metrics FID and RA from Table 1,  3, and  5. In Section  B.3, we’ve provided more detailed experiments to examine the necessity of decomposition in modeling high-frequency components with semantics, through which we have determined that the performance of frequency decomposition significantly surpasses that of merely removing too high-frequency noise.
@@
Method CRR(%, $\uparrow$) FID($\downarrow$) RA (%, $\uparrow$)
Ground-truth 13.02 0.03 97.78
OMOMO [42] 8.81 6.19 68.89
CG-HOI [20] 8.64 5.50 70.00
Ours 10.29 4.45 81.11
Table 4: Results on BEHAVE [2] dataset. The best in each column is highlighted in bold.
Figure 9: Qualitative results from TACO [48] dataset. Periodic relative motions are required between two objects. The color changes from deep to light, representing time passage. After removing the decomposition mechanism, the spatula tends to get stuck in a small area on the plate’s surface, without effective relative movements.
@@
Figure 10: User study results on different dataset splits.
Removal of $\mathcal{L}_{\text{align}}$ or Explicit Synchronization. As is shown in Figure 7 and 6, removing either one of $\mathcal{L}_{\text{align}}$ or explicit synchronization leads to unreasonable penetration or contact loss between objects or between humans and objects, with explicit synchronization playing a more significant role. This phenomenon is also revealed in quantitative evaluation results from Table  1,  3 and  5. An observed phenomenon is that simply incorporating frequency decomposition (w/o $\mathcal{L}_{\text{align}}$, exp sync) poses higher demand on synchronization, which is manifested as worse contact-based metrics (Defined in Section  4.2) than “w/o all”. Simply integrating $\mathcal{L}_{\text{align}}$ (w/o exp sync) can not fully solve this issue, making explicit synchronization steps become a must.
4.5 User Study
In order to evaluate semantic correctness of synthesized interactions from a perceptual aspect, we conduct a user study, comparing our method with all other baselines on the five datasets.
There are $10$ different splits in our experiments (TACO splits 1-4, CORE4D splits 1-2, OAKINK2 test split, GRAB splits of unseen subjects/objects, BEHAVE test split). We randomly draw 15 samples per split, resulting in a total of $150$ different questions ($15\times 10$). Each question requires participants to choose the best option among Ours and all other baselines, based primarily on “Completion, Realism, and Naturalness”. If the decision is difficult to make, participants are also instructed to consider “additional misalignments such as interpenetration, contact loss, jitter”. $150$ different individuals are involved, and each is assigned $10$ different questions, where the option orders are randomly shuffled. Our mechanism guarantees that each question is answered by exactly $10$ different agents. Figure  10 shows the times each method voted as the “best” in each split.
From the data of User Study (which is also corroborated by other metrics such as contact-based metrics and metrics regarding semantics), it can be observed that for datasets with fewer bodies, like GRAB or BEHAVE, the gap between our method and the baselines is not as significant as in datasets with more bodies. This is because, as the number of bodies increases, the requirements for synchronization become higher, and the interaction patterns between rigid objects become more complex, necessitating explicit modeling of high-frequency components through frequency decomposition, as well as those synchronization mechanisms. This further highlights the necessity of our approach in the general task of synchronized multi-body interaction synthesis.
5 Conclusions and Discussions
This paper presents SyncDiff, a unified framework for synchronized motion synthesis of multi-body HOI interaction by estimating both data sample scores and alignment scores, and jointly optimizing sample and alignment likelihoods in inference. We also introduce a frequency-domain decomposition to better capture high-frequency motions with semantics. Experiments on five datasets demonstrate that SyncDiff can be adapted to multiple scenarios with any number of humans, hands, and rigid objects. Comparative experiments also demonstrate that in each specific setting, our method achieves better contact accuracy and action semantics than a range of state-of-the-art baselines. Limitations, potential solutions, extensions, and some discussions of SyncDiff are provided in Section  E.
@@ A.1 Diffusion Model Basics
Diffusion models [75] and its variants [78, 30], especially latent diffusion, have been widely applied in different tasks, such as high-resolution image generation [67] or video generation [32]. They simulate the data distribution by introducing a series of variables $\{x_{i}\}_{i=1}^{T}$ with different noise levels. The forward noise process can be represented as
@@
For the inference process, from $T$ that is large enough so that $p(x_{T})\approx\mathcal{N}(0,I)$, do reverse sampling step by step, following
@@ A.2 Complete Proof of Our Two Synchronization Mechanisms
The goal in this section consists of two aspects:
1) Illustrate how to define a set of alignment scores featuring synchronization analogous to the commonly used data sample scores in diffusion models, and derive the corresponding loss term $\mathcal{L}_{\text{align}}$ in Section  3.5;
2) Prove that the explicit synchronization formulas in Section  3.6 are equivalent to maximum total likelihood sampling on the newly computed Gaussian distribution, where data sample scores and alignment scores are jointly considered.
Before we start, let’s derive some commonly used formulas in diffusion models that we will need in our proof, which the readers might not be familiar with. If you are familiar with the step-by-step denoising formula of diffusion, you can skip directly to Eq  12 and  13.
@@
In general, diffusion models predict the noise $\epsilon$, but in our setting, we need to define the alignment loss later. Thus, it’s more convenient to direct predict $\hat{x}_{0}$, which is the result after complete denoising. Note that
@@
By removing the coefficient, it becomes our simple reconstruction loss. Note that in our implementation, we split it into $\mathcal{L}_{\text{dc}}$ and $\mathcal{L}_{\text{ac}}$ for separate supervision. Similarly, in order to derive the alignment loss, our approach is still to express the loss function as some form of negative log-likelihood. Consider a triplet $(\hat{x}_{\textbf{b}_{1}},\hat{x}_{\textbf{b}_{2}},\hat{x}_{\textbf{b}_{2}\rightarrow\textbf{b}_{1}})$, where $\textbf{b}_{1}$, $\textbf{b}_{2}$ are two different bodies. We know that the definition of the score function is $\nabla\log p(x)$, which is just the negative gradient of the negative log-likelihood. As stated in the introduction part, we hope that this score function can guide $\hat{x}_{\textbf{b}_{2}\rightarrow\textbf{b}_{1}}$ towards $\text{rel}(\hat{x}_{\textbf{b}_{1}},\hat{x}_{\textbf{b}_{2}})$. A simple solution is to let $\hat{x}_{\textbf{b}_{2}\rightarrow\textbf{b}_{1}}$ follow a distribution $\mathcal{N}(\mu,\sigma^{2})$, where $\mu=\text{rel}(\hat{x}_{\textbf{b}_{1}},\hat{x}_{\textbf{b}_{2}})$, and $\sigma$ is a parameter that we can tune. Note that a vector $\hat{x}$ satisfying $\hat{x}\sim\mathcal{N}(\hat{\mu},\sigma^{2}I)$ has the probability density function
@@
It is not difficult to obtain that the negative log-likelihood is $\frac{1}{2\sigma^{2}}\|\hat{x}_{\textbf{b}_{2}\rightarrow\textbf{b}_{1}}-\text{rel}(\hat{x}_{\textbf{b}_{1}},\hat{x}_{\textbf{b}_{2}})\|_{2}^{2}$. Summing up for all such pairs $(\textbf{b}_{1},\textbf{b}_{2})$, and removing the coefficients, we can derive the alignment loss
$\mathcal{L}_{\text{align}}=$ $\sum\limits_{j_{1},j_{2}\in[1,m],j_{1}\neq j_{2}}\|\hat{x}_{o_{j_{2}}\rightarrow o_{j_{1}}}-\text{rel}(\hat{x}_{o_{j_{1}}},\hat{x}_{o_{j_{2}}})\|_{2}^{2}$ (17)
$+\sum\limits_{i\in[1,n],j\in[1,m]}\|\hat{x}_{h_{i}\rightarrow o_{j}}-\text{rel}(\hat{x}_{o_{j}},\hat{x}_{h_{i}})\|_{2}^{2}.$
Next, let’s prove the equivalence between the explicit synchronization formulas and maximum total likelihood sampling in inference. In our task, during the inference process, suppose we want to derive $\hat{x}^{{}^{\prime}}=\hat{x}_{t-1}$ from $\hat{x}=\hat{x}_{t}$, we first use the denoising backbone to predict the mean value of distribution $\hat{\mu}=\hat{\mu}_{t}$.
In our task, $\hat{x}^{{}^{\prime}}=[\hat{x}^{{}^{\prime}}_{o_{1}},\hat{x}^{{}^{\prime}}_{o_{2}},\dots,\hat{x}^{{}^{\prime}}_{o_{m}},\hat{x}^{{}^{\prime}}_{h_{1}},\hat{x}^{{}^{\prime}}_{h_{2}},\dots,\hat{x}^{{}^{\prime}}_{h_{n}},\\ \hat{x}^{{}^{\prime}}_{o_{1}\rightarrow o_{2}},\hat{x}^{{}^{\prime}}_{o_{1}\rightarrow o_{3}},\dots,\hat{x}^{{}^{\prime}}_{o_{m}\rightarrow o_{m-1}},\hat{x}^{{}^{\prime}}_{h_{1}\rightarrow o_{1}},\hat{x}^{{}^{\prime}}_{h_{1}\rightarrow o_{2}},\dots,\\ \hat{x}^{{}^{\prime}}_{h_{n}\rightarrow o_{m}}]$, and $\hat{\mu}=[\hat{\mu}_{o_{1}},\hat{\mu}_{o_{2}},\dots,\hat{\mu}_{o_{m}},\hat{\mu}_{h_{1}},\hat{\mu}_{h_{2}},\dots,\hat{\mu}_{h_{n}},\\ \hat{\mu}_{o_{1}\rightarrow o_{2}},\hat{\mu}_{o_{1}\rightarrow o_{3}},\dots,\hat{\mu}_{o_{m}\rightarrow o_{m-1}},\hat{\mu}_{h_{1}\rightarrow o_{1}},\hat{\mu}_{h_{1}\rightarrow o_{2}},\dots,\\ \hat{\mu}_{h_{n}\rightarrow o_{m}}]$, so
$\frac{1}{2\sigma^{2}}\|\hat{x}^{{}^{\prime}}-\hat{\mu}\|_{2}^{2}$ $=\frac{1}{2\sigma^{2}}\left(\sum\limits_{i=1}^{n}\|\hat{x}^{{}^{\prime}}_{h_{i}}-\hat{\mu}_{h_{i}}\|_{2}^{2}\right.$
$\quad+\sum\limits_{j=1}^{m}\|\hat{x}^{{}^{\prime}}_{o_{j}}-\hat{\mu}_{o_{j}}\|_{2}^{2}$
$\quad+\sum\limits_{j_{1},j_{2}\in[1,m],j_{1}\neq j_{2}}\|\hat{x}^{{}^{\prime}}_{o_{j_{2}}\rightarrow o_{j_{1}}}-\hat{\mu}_{o_{j_{2}}\rightarrow o_{j_{1}}}\|_{2}^{2}$
$\quad+\left.\sum\limits_{i\in[1,n],j\in[1,m]}\|\hat{x}^{{}^{\prime}}_{h_{i}\rightarrow o_{j}}-\hat{\mu}_{h_{i}\rightarrow o_{j}}\|_{2}^{2}\right)$
@@
$\mathcal{P}_{\text{align}}$ $=\sum\limits_{v=1}^{|V|}\lambda_{v}\|c_{v}-a_{v}\circ b_{v}\|_{2}^{2}$